WLCG-OSG-EGEE Operations meeting

28-R-15 (CERN conferencing service (joining details below))


CERN conferencing service (joining details below)

Nick Thackray
Weekly OSG, EGEE, WLCG infrastructure coordination meeting.
We discuss the weekly running of the production grid infrastructure based on weekly reports from the attendees. The reported issues are discussed, assigned to the relevant teams, followed up and escalated when needed. The meeting is also the forum for the sites to get a summary of the weekly WLCG activities and plans
  • OSG operations team
  • EGEE operations team
  • EGEE ROC managers
  • WLCG coordination representatives
  • WLCG Tier-1 representatives
  • other site representatives (optional)
  • GGUS representatives
  • VO representatives
  • To dial in to the conference:
    a. Dial +41227676000
    b. Enter access code 0148141

    OR click HERE
    (Please specify your name & affiliation in the web-interface)

    Click here for minutes of all meetings

    Click here for the List of Actions

    Recording of the meeting
      • 4:00 PM 4:00 PM
        Feedback on last meeting's minutes
      • 4:01 PM 4:30 PM
        EGEE Items 29m
        • <big> Grid-Operator-on-Duty handover </big>
          From: SouthEast Europe and DECH
          To: North Europe and CERN

          Report from SouthEast Europe::
          • During the week the GGUS Ticket #42172 was transfered to political instances. The last SAM are ok now, but it is still opened. We believe the problem is fixed.
          • The GGUS Ticket #40707(Political) is duplicated GGUS #42008. The second one needs more time to solve the problem and the first one was extended.

          Report from DECH:
          • Two tickets (42199, 42015) unanswered for one site, ITPA-LCG2, both transfered to the next step, political instance.
          • The situation with the ticket 40521 is not clear, the case is at the last step of escalation for long time - corresponding ROC should review the situation on operational meeting.
        • <big> PPS Report & Issues </big>
          Please find Issues from EGEE ROCs and general info in:

        • <big> gLite Release News</big>
        • <big> EGEE issues coming from ROC reports </big>
          • ROC NE: Vera Hansper has submitted GGUS ticket 42341 on October 15th. This ticket has been assigned the the GridView support unit. No action has been undertaken yet to solve this ticket.
          • ROC UK/I: Noticed some problems with the formatting of this ROC report. Multiple unlabelled lines for the downtime history against many sites and only some of these line up.
        • <big> Publishing "correct" values for cluster SI2k and Total CPU</big>
          Last week PIC asked:
          In the GridMap monitoring (http://gridmap.cern.ch) if one clicks the "show SI2k" button in the "topology view" section, the sites are scaled wrt the "total cpus" value in a SI2k units, which looks as computed just multiplying the number of job slots published times the GlueHostBenchmarkSI00. As most of the clusters are not homogeneous, this is not correct.
          GlueHostBenchmarkSI00 is just the value to which internal accounting is normalized.

          Answer: If you scale CPU speed to some value then you should also scale your SubCluster's Total and Logical CPU count so that your total power is reflected.
      • 4:30 PM 5:00 PM
        WLCG Items 30m
        • <big> Job submission mechanism needed in the CREAM CE</big>
          What submission mechanisms does the CREAM CE need to support?
          - GRAM (WS / pre-WS)?
          - Condor?
          - etc.
        • <big> Pool accounts on VO Boxes </big>
        • <big> WLCG issues coming from ROC reports </big>
          1. ROC UK/I: very few sites passing the LHCb SAM tests today.
          2. ROC CERN: [Information] Support for a new Role=pilot for LHCb has been deployed at CERN
          3. ROC CERN: [Information] The ATLAS grid queues default resource requirements have been updated to select nodes with at least 1.9GB of memory
        • <big>WLCG Service Interventions (with dates / times where known) </big>
          Link to CIC Portal (broadcasts/news), scheduled downtimes (GOCDB) and CERN IT Status Board

          Many interventions scheduled this week. Please consult the URLs above for details.
          • CERN: on the 28/10 ten of our CEs will be down for about 30min each because of a scheduled hardware intervention.

          Time at WLCG T0 and T1 sites.

        • <big> WLCG Operational Review </big>
          Speaker: Harry Renshall / Jamie Shiers
        • <big> Alice report </big>
          1. Alice has produced a document describing the installation and required set-up of the CREAM CE.
        • <big> Atlas report </big>
          1. Item
        • <big> CMS report </big>
          1. Item
          Speaker: Daniele Bonacorsi
        • <big> LHCb report </big>
          1. Item
        • <big> Storage services: Recommended base versions </big>
          The recommended baseline versions for the storage solutions can be found here: https://twiki.cern.ch/twiki/bin/view/LCG/GSSDCCRCBaseVersions

        • <big> Storage services: this week's updates </big>
      • 5:00 PM 5:30 PM
        OSG Items 30m
        Speaker: Rob Quick (OSG - Indiana University)
        • Discussion of open tickets for OSG
          List of open tickets
      • 5:30 PM 5:35 PM
        Review of action items 5m
      • 5:35 PM 5:35 PM